OpenStudy (anonymous):

-7y=-15x-42 -7y=-6x+21 solve for x and y

OpenStudy (4n1m0s1ty):

You can use elimination here: -7y = -15x-42 - -7y= -6x +21 ------------------ 0 = -9x - 63 63 = -9x x = -7 Plug back into one of the original eq. -7y = -15(-7) - 42 y = whatever that comes out to

OpenStudy (anonymous):

Subtract equation 2 from equation 1. You will get your answer :)

OpenStudy (anonymous):

is y 105

OpenStudy (anonymous):

try again

OpenStudy (anonymous):

is it -22

OpenStudy (anonymous):

try again... use the force... (since today is May Force)

OpenStudy (anonymous):

what do you do to get y

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

so the order pair (-7, -9) is the intersection of both those lines...